In lab, we reacted 4-phenylbutyric acid with methanesulfonic acid to generate α-tetralone. How would the calculations for the theoretical yield look? If I, for example, had 0.500g of 4-phenylbutyric acid, how would the equation look? I know it's an intramolecular reaction, but I'm confused on how to write out of the equation for the theoretical yield.

Typically it is: (mass of reactant) x (1 mol / molecular weight of reactant) x ( molar ratio of product / 1 mol of reactant) x (molecular weight of product / 1 mol) = theoretical yield.
Answer #1

In the picture, you can see the balanced chemical reaction that I have.

Balanced chemical reaction

Polyphosphoric acid

Methanesulfonic acid

C10H12O2 C10H12O + H2O

Molecular weight: 4-phenylbutyric acid = (10 x 12) + (12 x 1) + (2 x 16) = 164 g

  -tetralone = (12 x 10) + (10 x 1) + (16 x 1) = 146 g

Theoretical yield

164 g of C10H12O2   146 g of C10H12O

0.5 g of C10H12O2 x g of C10H12O

x = (0.5 x 146) / 164

x = 73 / 164

x = 0.44

Theoretical yield =  0.44 g of  C10H12O